OpenStudy (anonymous):

solve the equation. identify any extraneous solutions. square root of "a"=-2 A. 4 is a solution of the original equation B. 4 is a solution of the original equation. -4 is an extraneous solution C. -4 is a solution of the original equation. -4 is an extraneous solution D. no solution Simplify the radical expression by rationalizing the denominator. 5/ square root of 30 A. 30/ square root of 5 B. square root of 30 C. square root of 180/30 D. square root of 30/6

Nnesha (nnesha):

\[\huge\rm \sqrt{a} = -2\] do you know how to solve for a ?

OpenStudy (anonymous):

no

Nnesha (nnesha):

to cancel square root you have t o take square both sides square root can be written as 1/2 exponent so when you take square root \[\huge\rm \sqrt{x} = (x)^\frac{1}{2}\] \[\huge\rm \sqrt{x} = ((x)^\frac{1}{2})^2\] \[\huge\rm x^{\frac{1}{\cancel{2}} \times \cancel{2}}\] = x so that's why take square both side so you can cancel out square root

OpenStudy (anonymous):

oh

Nnesha (nnesha):

yeah so take square both side let me know what you get :-)

OpenStudy (anonymous):

I got b

Nnesha (nnesha):

\[(\sqrt{a})^2 = (-2)^2\] a = 4 extraneous solution when you plug in a value into the original equation if you get equal sides then 4 isn't extraneous solution if both sides are not equal then 4 is extraneous solution

Nnesha (nnesha):

a = 4 so \[\sqrt{4} = -2\] both sides are equal ?

Nnesha (nnesha):

LAG >.<